# Single Image Test-Time Adaptation for Segmentation

### v1

@@ Abstract
Test-Time Adaptation (TTA) methods improve the robustness of deep neural networks to domain shift on a variety of tasks such as image classification or segmentation. This work explores adapting segmentation models to a single unlabelled image with no other data available at test-time. In particular, this work focuses on adaptation by optimizing self-supervised losses at test-time. Multiple baselines based on different principles are evaluated under diverse conditions and a novel adversarial training is introduced for adaptation with mask refinement. Our additions to the baselines result in a 3.51 and 3.28 % increase over non-adapted baselines, without these improvements, the increase would be 1.7 and 2.16 % only. [^1] Project website with code: https://klarajanouskova.github.io/sitta-seg/
TTA
Test-Time Adaptation
TTT
Test-Time Training
MAE
Masked Autoencoders
SAM
SegmentAnything Model
Ent
Entropy-Minimization
dIoU
Deep-Intersection-over-Union
Ref
Mask Refinement
Adv
Adversarial-Attack
PL
Pseudo-Labelling
AugCo
Augmentation-Consistency
ACDC
Adverse Conditions Dataset with Correspondences
IoU
Intersection over Union
mIoU
mean Intersection over Union
CE
Cross-Entropy
ViT
Vision Transformer
BCE
Binary Cross Entropy
FGSM
Fast Gradient Sign Method
PGD
Projected Gradient Descent
@@ 1 Introduction
Models at inference time often process data statistically different from the training data if the distribution changes or the data are from a low-probability area of the distribution. Poor performance on such data can be expected, which has encouraged research on unsupervised Test-Time Adaptation (TTA) techniques, i.e., adapting the model at test-time without the training data or information about the distribution. Impressively, prior work [8, 16] showed that model performance can be improved even by adapting to a single unlabelled image in isolation.
This work explores single, isolated image test-time adaptation for segmentation, assuming no other data are available. This allows for a thorough analysis of the method’s performance on different kinds and levels of covariate shifts while eliminating other factors. No assumptions are made about the distribution of subsequent images or the number of images available, making the method very general and applicable even when the model receives independent images. An example of masks before and after test-time adaptation are shown in Figure 1.
image ground truth NA TTA
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 1: Non-adapted (NA) and adapted (TTA) masks.
The single image TTA is unlike most TTA [41, 44] for segmentation, which focuses on continual test-time adaptation in a changing environment, either by gradually updating the model parameters or by collecting image statistics to be used for adaptation to individual images. While practical in many applications, such as autonomous driving, it is challenging to understand the strengths and weaknesses of different methods since evaluation is typically done on long sequences of images of varying levels and kinds of domain shift. The methods are evaluated on different models and datasets in different settings and often only feature the entropy-minimization [42] and batch-normalization [13] statistics adaptation [35, 25] baselines. Meanwhile, many methods improving over these baselines have been proposed in recent years both for segmentation and classification [28] but it is not known how these methods compare in different settings. Methods on continual TTA focusing on reducing catastrophic forgetting for methods based on optimizing the model weights, such as [41] are often agnostic to the loss function and can be used with any of the single-image adaptation methods studied in this work, the improvements being orthogonal.
There is a line of work on TTA for segmentation in medical imaging [15, 40] that has been independent of the vision community. Still, the methods are general and can be applied to non-medical images as well. These methods are based on optimizing self-supervised loss functions with learned mask discriminators or mask refinement modules. Only reconstruction loss, which requires changes to the model architecture and training process, was evaluated on segmentation before, and self-supervised loss functions for image segmentation TTA remain underexplored.
Apart from classification methods adapted to segmentation and previous segmentation baselines, our work explores methods similar to those proposed for medical imaging. In particular, we focus on learning a mask-refinement module to produce enhanced pseudolabels at test-time. Novel adversarial training of the mask refinement module is introduced, replacing the originally proposed heuristic of swapping image patches. Swapping patches results in unrealistic images and requires finding a suitable size of the patches and the ratio of swapped patches. Controlling the severity of mask corruption is hard. This work simulates domain shift impact on masks by projected gradient descent adversarial attack on the input image, resulting in realistic masks in the first iterations of the attack. The attack learning rate and number of iterations then control the severity of the domain shift.
This work studies general methods that can be applied to existing pretrained segmentation models without assumptions of specific network layers or adaptations to the segmentation training process or model architectures. For this reason, methods based on image reconstruction [45] and updating batch normalization [26] layer statistics are not included in our experiments. Multiple methods novel to image segmentation are introduced, in particular Adversarial-Attack (Adv), Mask Refinement (Ref) and Deep-Intersection-over-Union (dIoU). The Augmentation-Consistency (AugCo) method is evaluated in the single image setup for the first time.
Experiments are conducted with two different pretrained segmentation models, one trained on the GTA5 [30] datasets and tested on Cityscapes [5] and ACDC [34] benchmarks. The other model is trained on COCO [20] and evaluated on VOC [6]. All hyper-parameter tuning and most method performance analyses are done on an extension of the training datasets with synthetic corruptions featuring a wide range of corruptions at different levels inspired by [12]. This has the advantage over synthetic datasets such as Synthia. Such a dataset can be obtained from an arbitrary segmentation training dataset. The main contributions of our work are:
@@
    We introduce image segmentation TTA baselines inspired by methods from other tasks and domains.
@@
    Novel adversarial refinement module training for Ref-based TTA
@@
    Improvements of baselines in single-image setup by replacing Cross-Entropy (CE) loss with Intersection over Union (IoU) loss.
@@
    Analysis of methods’ performance on different domain shifts and domain shift levels.
@@ 2 Related Work
Domain adaptation methods share a lot in common with semi-supervised learning [31, 33] and may exploit strategies such as generative modeling [39, 47]. In practice, this is often infeasible since source data may not be available or domain shift is not known beforehand. Various modifications of the domain adaptation scenario have recently emerged, for example by considering no access to source data or a continual domain shift [22, 41, 44, 3]. In particular, test-time adaptation methods assume that no source data is available. The differences between TTA and other setups are discussed in Appendix A.
Some works distinguish TTA and Test-Time Training (TTT), the difference being that TTA methods such as [27, 43] can be applied to arbitrary models without any additional constraints while TTT methods like [9, 3, 21] require modifications to the training process. However, the boundary is not clear and not all works make this distinction. This work refers to all methods as TTA.
There are two predominant approaches to test-time adaptation. The first approach adapts to input $x$ by minimizing a self-supervised loss function $\mathcal{L}_{\text{TTA}}$ over a set of parameters $\omega\in\theta$ of the segmentation network $f_{\theta}$: $\omega^{\star}=\operatorname*{arg\,min}_{\omega}\mathcal{L}_{\text{TTA}}(f,x)$.
The implementation of $\mathcal{L}_{\text{TTA}}(f,x)$ is what distinguishes the different methods. There are many possible choices for $\omega$ and the right choice may depend on time constraints, available GPU memory and the amount of data. In this work, two options are considered: $\omega=\theta$ where all the learnable parameters are optimized and $\omega=\theta_{n}$ where only the learnable parameters of the normalization layers such as BatchNorm, GroupNorm or LayerNorm are optimized. Details of $\mathcal{L}_{\text{TTA}}(f,x)$ implementation for the baseline methods and the necessary adaptations to the single-image setup are provided in Appendix B.
The second approach involves updating the mean and variance of the batch normalization layers as a weighted combination of the source and target statistics during the testing phase.
@@
Test-Time Adaptation methods for classification Many recent methods propose improved strategies to update the batch normalization statistics [35, 25]. A limitation of these methods is the reliance on presence of batch nromalization, which is often not part of recent transformer-based architectures. In [43], the learnable parameters of the normalization layers are also updated via entropy minimization. While this method is often reported as unstable since single-image statistics may not be sufficient, the method can also only update the normalization layers learnable parameters, without the statistics update, making it generalizable to all currently used architectures.
@@
On classification tasks, many methods outperforming the aforementioned baselines have been proposed. A combination of self-supervised contrastive learning to refine the features and online label refinement with a memory bank is proposed in [4]. Recently, a method based on updating the parameters of the normalization layers of the network by optimizing it for robustness against adversarial perturbation as a representative of domain shift was proposed in [27], outperforming similar test-time adaptation approaches. Rotation prediction is proposed in [37] as self-supervised task to be learnt alongside the main one and then optimized at inference time. Lately, it was shown that reconstruction with masked auto-encoders is a very strong self-supervised task for test-time adaptation of classifiers by [9].
@@
Test-Time Adaptation methods for segmentation To the best of our knowledge, the only work [16] also focused on adaptation to a single isolated image is based on computing the statistics from augmented version of the input image, assuming batch-normalization layers are present in the network. Both [29] and [44] exploit augmented views of the input images to identify reliable predictions. The method of [29] is based on the consistency of predictions between augmented views, which replaces prediction confidence for selecting reliable pixels. Cross entropy loss is then minimized on such reliable predictions, together with a regularization based on information entropy [19] to prevent trivial solutions. The method achieves impressive results, however, in contrast to our experiments, knowledge of the target domain shift is used for hyper-parameter tuning. The evaluation assumed a full test set available at once, focusing on source-free domain adaptation, rather than TTA, but the method is applicable to the TTA setup as well. In [41], the performance of entropy minimization in a continual setup is explored, proposing parameter restart to tackle weight drift, significantly improving performance. The focus is on driving datasets only. Similarly, [44] also focus on continual adaptation. Again, augmentations of the images are generated to obtain more reliable predictions. Further, the network parameters are stochastically reset to their initial values to prevent forgetting of the source domain knowledge.
@@
Test-Time Adaptation methods for medical imaging In [15], an autoencoder is proposed that translates predicted masks into refined mask. At test time, the segmenter is optimized to produce masks closer to the enhanced ones. However, this work assumes the whole test dataset is available at once, in contrast to our single-image setup. The work of [40] is similar to [15] but instead of a masked-autoencoder, a GAN-like discriminator trained end-to-end together with the segmenter is used, as well as an auxiliary reconstruction loss.

### v2

# Test-Time Adaptation TTT Test-Time Training

@@ Abstract
Test-Time Adaptation (TTA) methods improve domain shift robustness of deep neural networks. We explore the adaptation of segmentation models to a single unlabelled image with no other data available at test time. This allows individual sample performance analysis while excluding orthogonal factors such as weight restart strategies. We propose two new segmentation TTA methods and compare them to established baselines and recent state-of-the-art. The methods are first validated on synthetic domain shifts and then tested on real-world datasets. The analysis highlights that simple modifications such as the choice of the loss function can greatly improve the performance of standard baselines and that different methods and hyper-parameters are optimal for different kinds of domain shift, hindering the development of fully general methods applicable in situations where no prior knowledge about the domain shift is assumed.
Code and data: https://klarajanouskova.github.io/sitta-seg/
@@ 1 Introduction
A common challenge in machine learning stems from the disparity between source (training) and target (deployment) data domains. Models optimized to minimize an error on a dataset from a specific domain are often expected to perform reliably in different domains. The discrepancy between training and deployment data, known as the domain shift, is very common; in fact, few things do not change in time, and training happens (well) before deployment. A domain shift may substantially degrade model performance at deployment time despite proper validation on training data, yet it is often not explicitly addressed and most machine learning effort has focused on the generalization problem.
In many practical scenarios, the characteristics of the target domain are not known beforehand, making the preparation of the model with traditional domain adaptation Rodriguez & Mikolajczyk (2019); Tzeng et al. (2017) techniques nontrivial. Recent advances Wang et al. (2020b); Sun et al. (2020); Gandelsman et al. (2022) suggest that under certain weak assumptions about the domain shift - such as a stable label distribution across domains - it is possible to mitigate the performance degradation with methods based on the information carried by input data received at inference time.
Test-Time Adaptation (TTA) is suitable for a priori unknown or difficult-to-predict domain shifts. Characterized as an unsupervised and source-free technique, TTA adapts the model directly during inference. The source-free nature, i.e. without access to the original training data, ensures compliance with data governance standards and enables adaptation in memory-constrained environments.
Single Image Test-Time Adaptation (SITTA) tailors a model at test time to each individual image. Since it operates on a single image, it does not introduce assumptions about the stability of the data distribution over time. Each time starting from the weights fixed at training time, SITTA is safe to use when any form of memorization of the deployment data is prohibited. A disadvantage is an increased computational time compared to batch methods and the lack of retaining the acquired knowledge. On the other hand, SITTA could be leveraged to reduce the computational cost of adapting to a sequence of similar images by only processing the most informative samples. Despite the advantages, only Khurana et al. (2021) primarily addresses SITTA, the mainstream of TTA research deals with continual test-time adaptation in a changing environment Niu et al. (2023); Volpi et al. (2022); Wang et al. (2022). These methods typically gradually update model parameters or accumulate image statistics for subsequent adaptation to individual images. Continual TTA strategies are practical in many applications, such as autonomous driving, but they are challenging from the point of view of accurate assessment of relative strengths and weaknesses. The difficulty arises from the evaluations conducted on many images with varying levels and kinds of domain shift. Moreover, the sequence in which images are presented can significantly influence performance metrics, adding a layer of complexity to assessing the true efficacy of these methods. SITTA streamlines the evaluation process and contributes to understanding the broader class of TTA strategies. Continual TTA analysis has focused on complementary issues such as catastrophic forgetting.
Figure 1: The proposed experimental framework for Single Image Test-Time Adaptation (SITTA). Hyper-parameters are found on a synthetic dataset derived from the training set by applying a diverse set of corruptions. SITTA methods are then tested on real-world datasets with domain shift.
In this paper, we explore and improve the state of the art in Single Image Test-Time Adaptation for image semantic segmentation. While TTA has been applied to many tasks; the overwhelming majority of methods have been developed for image classification. Image segmentation TTA methods are each evaluated under very different conditions and compared to a limited set of baselines, making understanding their performance difficult. We focus on semantic segmentation SITTA with self-supervised loss functions. These methods are broadly applicable across various segmentation models and domains without being constrained by specific architectural or segmentation training requirements. Consequently, the many works that rely on batch normalization layers are not included in this work as they are incompatible [^1] At inference time, batch normalization uses per-feature mean and variance computed over the training dataset. TTA methods also incorporate the test data feature statistics. Transformer models typically use layer-normalization layers, which normalize features for each instance independently and thus always use the test sample statistics. with the modern transformer architectures currently dominating the field. Likewise, methods such as image-reconstruction based TTA Wang et al. ; Gandelsman et al. (2022) are not included due to their significant demands for training process modification and model architecture adjustments. Some methods require access to training data before deployment since they train an auxiliary network. However, they do not alter the training of the segmentation model.
Six TTA methods are evaluated in the SITTA setting. Two are established baselines: entropy minimization Hu et al. (2019) and pseudolabelling Lee et al. (2013) (self-training). They are the only ones evaluated in the SITTA setup before. The third method Prabhu et al. (2021) is a recently proposed segmentation TTA method that identifies confidently pseudolabelled pixels as consistent across augmentations. The fourth method Nguyen et al. (b) was proposed for image classification; we are the first to use it for image segmentation. The main idea is that if a network is robust to small adversarial domain shifts (it is optimized to not change the prediction under an adversarial attack at test time), it will also be robust to real-world domain shifts.
The remaining two approaches are novel. We combine and extend two techniques optimizing self-supervised losses with 1. learnt segmentation mask domain discriminators Tzeng et al. (2017), so far used for unsupervised domain adaptation and never applied to TTA, and 2. Segmentation mask refinement  [^2] Called “denoising autoencoders” in Karani et al. (2021). modules Karani et al. (2021); Valvano et al. (2021), never considered outside the medical domain TTA.
The idea of training-time unsupervised adaptation [^3] In the domain adaptation literature, the term “unsupervised” refers to the situation when images from the target domain are available to the method, but they are not labelled.In such a setting, training a source-target domain discriminator is possible, which contradicts the standard implication of the term “unsupervised”. The setting where no samples from the target domain are available during training is referred to as “test-time adaptation” or “unsupervised source-free adaption”. with mask discriminators is that a classifier trained to distinguish masks from the source and target domains can be used to supervise adaptation. The source-domain segmentation model is finetuned to produce masks on the target domain that resemble those of the source domain, as predicted by the discriminator. This method is not directly applicable to TTA since the target images to train the discriminator are unavailable. However, the target domain masks can be synthesized. Copying random patches into different image locations was proposed in the medical domain. Such augmentation makes sense when the textures are more important than the shapes of objects, such as in the medical domain, but are not general enough for other images. We propose to use the output of the segmenter on adversarially attacked images. The adversarial optimization learning rate and the number of iterations control the severity of the domain shift. Further, Karani et al. (2021) proposes to replace the mask discriminator with a denoising autoencoder. Since a discriminator can use shortcuts and only focus on the most discriminative parts of the masks, a model that produces a refined mask is trained to consider all the masks. We explore both options.
We address the problem that arises from the practice in the current landscape of segmentation TTA that the performance assessment is carried out with inconsistent adaptation settings. For instance, keeping batch normalization statistics constant or updating them during entropy minimization can yield substantially different outcomes. If only one of the two options is tested Khurana et al. (2021); Volpi et al. (2022), contradictory results are reported. Often, the evaluation compares only with the established baselines such as entropy minimization Wang et al. (2020a) and batch normalization Ioffe & Szegedy (2015a) statistics adaptation Schneider et al. (2020); Nado et al. (2020), ignoring recent progress. Many methods that improve these baselines have been proposed in recent years for both segmentation and classification Nguyen et al. (a); Gao et al. (2023); Niu et al. (2023); Chen et al. (2022), and their relative merit is unknown, since a comprehensive comparison with a well-defined methodology is lacking.
Our experimental framework, depicted in Figure 1, consists of three stages: 1. Segmentation model training, which is standard. 2. Test-time adaptation training which tunes TTA hyper-parameters. We utilize an augmented version of the training datasets. This extension incorporates synthetic corruptions inspired by Hendrycks & Dietterich (2019). The corruption types include different kinds of noise and blur, weather conditions such as fog or frost, jpeg compression, and basic image intensity transformations. It can be derived from an arbitrary segmentation training dataset, as opposed to existing synthetic datasets used by previous work such as Ros et al. (2016). It provides precise control over the conditions and facilitates detailed analysis. 3. Evaluation on real-world test datasets with domain shift.
We experiment with two existing pretrained models for semantic segmentation. The models have different architectures and are trained on various datasets. Almost all segmentation TTA methods are evaluated on driving scenes benchmarks - we follow this practice and add a benchmark on common objects.
The main contributions of this paper are:
@@
    We conduct a comparative study of six TTA techniques run in SITTA mode for image segmentation: Two established baselines, two adapted state-of-the-art methods from image classification and continual TTA, and two proposed methods.
@@
    Novel methods adapting ideas from unsupervised domain adaptation and medical imaging TTA to non-medical image segmentation are introduced, filling a gap in exploring diverse self-supervised loss functions. The method outperforms the other methods on multiple test datasets and is shown to be powerful on the images with the worst segmentation performance, as measured by Intersection over Union (IoU).
@@
    Improvements of baselines in the single-image setup by replacing Cross-Entropy (CE) with the IoU loss. The performance of pseudo-labelling is improved by 3.51 % and 3.28 % on GTA5-C and COCO-C validation sets while with CE loss, the improvements are by 1.7 % and 2.16 % only, respectively.
@@
    The first work shows the potential of SITTA for segmentation, an underexplored setup essential in applications with strict data governance standards or high variability among individual images.
2 Background
Common approaches to domain adaptation change the style of labelled source images to resemble the training images Tzeng et al. (2017); Zhang et al. (2018) or train domain classifiers to guide the adaptation process. In practice, this is not always feasible since source data may not be available for example for privacy or memory limitation reasons, or we may only have a small number of target domain images available when data arrive individually/in small batches, rather than all at once. In continually evolving environments, the distribution may change by the time adaptation on a large target dataset is completed. Various modifications of the traditional domain adaptation scenario tackling the aforementioned limitations have recently emerged, for example by considering no access to source data or a continual domain shift Liu et al. (2021a); Volpi et al. (2022); Wang et al. (2022); Bartler et al. (2022).
In particular, test-time adaptation methods assume no source data is available and aim to exploit the information from as little as a single target domain image. Like other domain adaptation methods, TTA methods are often inspired by semi-supervised learning methods. For instance, the most common TTA baseline relies on minimization of the predictions entropy, a method inspired by Saito et al. (2019). Other methods rely on adapting the batch normalization statistics, inspired by methods like adaptive batch normalization Li et al. (2018), or aggregating statistics to create so-called prototypes Tanwisuth et al. (2021) that can be used to build a classifier.
Some works also distinguish between TTA and Test-Time Training (TTT). The difference between TTA and TTT is that TTA methods such as Nguyen et al. (b); Karani et al. (2021) can be applied to arbitrary pre-trained models without any additional constraints while TTT methods like Gandelsman et al. (2022); Bartler et al. (2022); Liu et al. (2021b) require modifications to the training process. However, not all works make this distinction and the boundary is not always clear, as some methods like Karani et al. (2021) need to train an auxiliary deep net on the source data but do not modify the model pretrained weights. In this work, both will be jointly referred to as TTA for simplicity.
In Appendix A, other related domain adaptation scenarios and their relation to TTA are described.
Generally, TTA methods can be split into three groups: Adaptation in the input space, feature space and output space. Input space adaptation aims to translate the images from the source domain to the input domain. In practice, Gao et al. (2023) achieve this by feeding target images with added noise to a diffusion model trained on the source data, coupled with reconstruction guidance to preserve semantics. The model doesn’t retain any knowledge from the adaptation - the advantage is it is not susceptible to catastrophic forgetting but the disadvantage is it may limit the adaptation capabilities. Adaptation in the feature space is the most common approach and typically relies on optimizing the network parameters via a self-supervised loss function. This can be done directly, i.e. through prediction entropy minimization, or by training an auxiliary task such as image reconstruction. Another set of feature-adaptation approaches are parameter-free and rely on accumulating the image statistics, such as the mean and variance of image features, or by aggregating confident prediction features into so-called prototypes, which are then used for classification. Output space adaptation techniques aim to improve the network output without neither altering the network parameters and statistics nor the input image. This is done for instance in Karani et al. (2021) where an auxiliary network is trained to predict a refined mask. To the best of our knowledge, output space adaptation methods are typically only used to provide pseudo-masks, turning them into feature-space adaptation methods. This helps to iteratively improve the pseudo-masks and adapt to larger domain shifts. All the methods evaluated in this work can be considered as feature space adaptation methods, possibly via output space adaptation.
@@ 3 Related Work
Test-Time Adaptation methods for classification. Many recent methods propose improved strategies to update the batch normalization statistics Schneider et al. (2020); Nado et al. (2020). A limitation of these methods is the reliance on presence of batch nromalization, which is often not part of recent transformer-based architectures. In Wang et al. (2020b), the learnable parameters of the normalization layers are also updated via entropy minimization. While this method is often reported as unstable since single-image statistics may not be sufficient, the method can also only update the normalization layers learnable parameters, without the statistics update, making it generalizable to all currently used architectures.
@@
On classification tasks, many methods outperforming the aforementioned baselines have been proposed. A combination of self-supervised contrastive learning to refine the features and online label refinement with a memory bank is proposed in Chen et al. (2022). Recently, a method based on updating the parameters of the normalization layers of the network by optimizing it for robustness against adversarial perturbation as a representative of domain shift was proposed in Nguyen et al. (b), outperforming similar test-time adaptation approaches. Rotation prediction is proposed in Sun et al. (2020) as self-supervised task to be learnt alongside the main one and then optimized at inference time. Lately, it was shown that reconstruction with masked auto-encoders is a very strong self-supervised task for test-time adaptation of classifiers by Gandelsman et al. (2022).
@@
Test-Time Adaptation methods for segmentation. To the best of our knowledge, the only work also focused on adaptation to a single isolated image Khurana et al. (2021) is based on computing the statistics from augmented version of the input image, assuming batch normalization layers are present in the network. Both Prabhu et al. (2021) and Wang et al. (2022) exploit augmented views of the input images to identify reliable predictions. The method of Prabhu et al. (2021) is based on the consistency of predictions between augmented views, which replaces prediction confidence for selecting reliable pixels. Cross entropy loss is then minimized on such reliable predictions, together with a regularization based on information entropy Li et al. (2020) to prevent trivial solutions. The method achieves impressive results, however, in contrast to our experiments, knowledge of the target domain shift is used for hyper-parameter tuning. The evaluation assumed a full test set available at once, focusing on source-free domain adaptation, rather than TTA, but the method is applicable to the TTA setup as well. In Volpi et al. (2022), the performance of entropy minimization in a continual setup is explored, proposing parameter restart to tackle weight drift, significantly improving performance. The focus is on driving datasets only. Similarly, Wang et al. (2022) also focus on continual adaptation. Again, augmentations of the images are generated to obtain more reliable predictions. Further, the network parameters are stochastically reset to their initial values to prevent forgetting of the source domain knowledge.
@@
Test-Time Adaptation methods for medical imaging. In Karani et al. (2021), an autoencoder is proposed that translates predicted masks into refined mask. At test time, the segmenter is optimized to produce masks closer to the enhanced ones. However, this work assumes the whole test dataset is available at once, in contrast to our single-image setup. The work of Valvano et al. (2021) is similar to Karani et al. (2021) but instead of a masked-autoencoder, a GAN-like discriminator trained end-to-end together with the segmenter is used, as well as an auxiliary reconstruction loss.
@@
Enhancing existing TTA benchmarks. There are multiple concurrent works that identify similar issues and reporting results consistent with our experiments, mostly for image classification. The work of Yu et al. (2023) also highlights the issue of evaluating each method under very different conditions and provides a benchmark for image classification TTA encompassing different adaptation scenarions, as well as diverse backbones and domain shift datasets. Similarly to ours, a significant disparity between synthetic corruptions performance and natural shifts is observed. However, the hyper-parameters were selected based on a single kind of domain shift, which may bias the results. Another work adressing the issue of fair comparison of TTA methods is that of Mounsaveng et al. (2024) which provides an analysis of existing orthogonal classification TTA methods. Class rebalancing is one of the tricks proposed to improve the methods’ performance. Also, sample filtration to remove noisy high-entropy images is employed. In contrast, we analyze performance of different methods based on prediction entropy, showcasing some methods can actually be highly effective on those noisy, high-entropy samples. Similarly to ours, the work shows that baselines can be greatly improved by very simple tricks. In Niu et al. (2023), label imbalance at test-time is again identified as an important factor harming the TTA performance. Again, the works focus is on image classification and epxlores different normalization layer kinds and stabilization techniques of entropy minimization while we focus on comparions of cross-entropy and a class-imbalance aware segmentation loss function, the IoU. Finally, Yi et al. (2023) study TTA for image segmentations and how well classification methods transfer to semantic segmentation TTA. They conclude that many of the classification TTA improvements do not transfer to segmentation and again highlight the class imbalance, which is typically greater for segmentation datasets.